rolling a die and getting an outcome that is less than 5
Answer #1

Solution:

The probability of getting any of the 6 possible outcomes (ie; 1,2,3,4,5,6) when a die is rolled is 1/6 , if the die is fair.

Now, for getting an outcome less than 5, any of the possible outcomes 1, 2, 3, or 4 will be favourable.

Since each of them have probability as 1/6, therefore the total probability will be 4*(1/6) = 4/6 = 2/3.